OpenStudy (alibaby):

Find a rational number that is between 9.5 and 9.7. Explain why it is rational.

OpenStudy (alibaby):

@rishavraj

OpenStudy (alibaby):

@Will.H @zepdrix

zepdrix (zepdrix):

Can you think of a decimal number between 9.5 and 9.7?

OpenStudy (alibaby):

9.6

zepdrix (zepdrix):

Ok good. But how do we know that this decimal value is rational? Well, we have a couple of options. We can simply say that it has a terminating decimal, therefore it must be a rational number. Or, and this is probably what they want us to do... we can write this decimal number as a ratio of whole numbers. That's what defines a rational number. For example: 3.5 is rational because it's the fraction 7/2. Two `whole numbers` being divided.

zepdrix (zepdrix):

Well we could write our number as \(\large\rm \frac{9.6}{1}\) but this is not a whole number on top. We can fix that by multiplying the top and bottom by 10.\[\large\rm \frac{9.6}{1}=\frac{96}{10}\]Yay! We have a whole number (integer) over another whole number, so it's rational. Yay team!

OpenStudy (alibaby):

thanks @zepdrix skype me